OpenStudy (anonymous):

lim x-> 0+ sin^-1 2x / x help please

OpenStudy (watchmath):

Are you asking \[\lim_{x\to 0^+}\frac{\sin^{-1} (2x)}{x}\]

OpenStudy (solomonzelman):

\(\Large\color{black}{\displaystyle\lim_{x \rightarrow ~0^+}~\frac{\sin^{-1}(2x)}{x}}\) ~ \(\Large\color{black}{\sin^{-1}(0)=0}\) sounds like it is an L'H'S problem.

hartnn (hartnn):

or plug in x = (sin y)/2 :)

hartnn (hartnn):

when x->0+ y->0+

OpenStudy (watchmath):

Let \(f(x)=\sin^{-1}(2x)\). Then the limit that you are looking for is the definition of \(f'(0)\).